write the equation of the line that passes through (0, 5(1/2)) and (-3(1/2), 0)
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

write the equation of the line that passes through (0, 5(1/2)) and (-3(1/2), 0)
---------
slope:: (5/2)/(-3/2) = -5/3
---------
Form::: y = mx + b
Solve for "b"::::
-------
(5/2) = (-5/3)(-3/2) + b
----
5/2 = -5/2 + b
b = 5
======
Equation:: y = (-5/3)x + 5
=============================
Cheers,
Stan H.
==============

Answer by mananth(16946)

x1 y1 x2 y2
0 5 1/ 2 -3 1/ 2 0

slope m = (y2-y1)/(x2-x1)
( 0 - 5 1/2 )/( -3 1/2 - 0 )
( -5 1/2 / -3 1/2 )
m= 1 4/ 7
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( 0 , 11/2 ) in
Y = m x + b
5.50 = 11/7 * 0 + b
5.50 = 0 + b
b= 11/2 + -0
b= 11/ 2
So the equation will be
Y = 11/7 x + 11/2
m.ananth@hotmail.ca